### v1

@@ 6 Cross-correlation Signal Dependency on GW Source and Lens Parameters
Figure 9: The figure shows the variation of the residual cross-correlation with respect to variation in the match-filtering SNR of the GW signal by varying the chirp mass (detector frame) and lensing parameters (b, $f_{0}$ and k). The cross-correlation timescale is chosen to be $\tau=0.5s$.
@@
The residual cross-correlation $(D_{xx^{\prime}})$ depends on the performance of the unlensed hypothesis parameter estimation to obtain the best-fit GW strain. The better the strain data matches with the best-fit strain is, the smaller is the residual and the residual cross-correlation. However, it is shown in the section 7 that, GW source parameters are well constrained with a model considering no lensing hypothesis even when there is microlensing present. This happens because the source parameters are completely non-degenerate with the lensing template considered in the analysis. Therefore, the residual cross-correlation depends on the strength of the signal, formally given by the match-filtering SNR, that defines how well can we obtain the best-fit. Thus since the microlensing parameters are not captured by the lens parameters in the unlensed hypothesis PE, there exists strongly dependency of the residual cross-correlation on the values of the microlensing parameters. Therefore, to observe how well can we distinguish a lensing modulation, we have checked the variation of the residual cross-correlation with respect to lensing parameters. In figure 9, we have shown the detectability of the residual cross-correlation with respect to the background noise cross-correlation, when no signal is present in the data.
@@
The shaded regions in figure 9, shows the 1-$\sigma$ bounds on the noise cross-correlation distribution. The lines show the residual cross-correlation value when the signal consists of microlensing amplification. We varied the chirp mass to increase the strength of the signal. We fixed the value of $a$ to 1.2 and 1.8, and then varied the parameters of $b$, $f_{0}$ and $k$. The luminosity distance of the source is kept fixed at $1$ Gpc. The values of the microlensing parameters are chosen in consistency with the numbers obtained in 3 when we fitted the amplification models of BH and SIS with our template in figure 4.
@@
In the first row of the plots, we have varied the value of the microlensing amplitude parameter $b$, keeping $f_{0}$ and $k$ fixed at $30$ Hz and $0.001$ respectively. We observe that for $a=1.8$, $b=0.4$ at chirp mass as low as $20M_{\odot}$ the cross-correlation signal at H-L lies 3-$\sigma$ above, however at L-V pair the cross-correlation signal comes above the noise distribution 3-$\sigma$ at chirp mass $40M_{\odot}$, same for V-H is at around $25M_{\odot}$. As the value of b is decreased to a lower value, the minimum chirp mass at which the cross-correlation signal is well distinguishable  [^10] By ”well distinguishable” we mean that, the residual cross-correlation is at least 3-$\sigma$ away from the noise cross-correlation distribution of the BBH shifts towards higher masses. Similarly for $a=1.2$ with $b=0.4$, $f_{0}=30$ Hz, $k=0.001$ we observe that for H-L, L-V, and V-H the cross-correlation signal is distinguishable at $30M_{\odot}$, $50M_{\odot}$ and $35M_{\odot}$ respectively, which is moved towards higher values as the value of b is decreased.
@@
In the second row of the plots, we have varied the value of the microlensing oscillation parameter $f_{0}$, keeping $b$ and $k$ fixed at $0.4$ and $0.001$ respectively. Unlike the cross-correlation dependency on $b$, we don’t observe a direct observational effect of $f_{0}$ variation, for any of the detector pairs. For both the cases with $a=1.8$ and $a=1.2$, we observe that the three lines corresponding to $f_{0}=30$ Hz, $f_{0}=60$ Hz, $f_{0}=100$ Hz, the lines are intertwined, not significantly altering the minimum chirp mass at which the cross-correlation signal is distinguishable. We understand that, the value of $b$ is more critical to determining microlensing in GW data as compared to $f_{0}$.
@@
In the third row of the plots, we have varied the value of the microlensing oscillation converging parameter $k$ (as the GW frequencies move towards strong lensing regime), keeping $b$ and $f_{0}$ fixed at $0.4$ and $30$ Hz respectively. Quite similar the cross-correlation dependency on $b$, we observe a strong dependency of $k$ variation. For the case with $a=1.8$, we can distinguish the cross-correlation signal for H-L, L-V, and V-H at $35M_{\odot}$, $45M_{\odot}$ and $40M_{\odot}$ respectively for the value of $k=0.008$. The minimum chirp mass is moved towards the lower end as the value of $k$ is decreased. Similarly, with $a=1.2$, we can distinguish the cross-correlation signal for H-L, L-V, and V-H at $40M_{\odot}$, $60M_{\odot}$ and $45M_{\odot}$ respectively for the value of $k=0.008$. We understand that the effect of the convergence parameter $k$ is similar to the effect of perturbation parameter $b$, therefore making it a very crucial parameter for the detection of microlensing signal.
7 Joint exploration of GW Source Parameters and Lensing Parameters
@@
Lensing can affect the inference of the source properties obtainable from the GW signal through existing parameter estimation techniques. Therefore, a correct approach for inferring source parameters, not just includes the source intrinsic and extrinsic parameters, rather it as well consists of the lens-induced characteristic parameters on the GW signal. We perform Bayesian parameter estimation technique in the frequency domain to jointly explore the parameter space occupied by the source and lens parameters. To showcase the capability of the technique, we choose a simple model for the lensing amplification factor
$A(f;a,b,k,f_{0})=a\left[1+be^{-kf}cos\left(2\pi\frac{f}{f_{0}}\right)\right],$ (7.1)
here, we have kept the value of $\phi=0$ for the simulation set for the purpose of demonstrating the method. The results including $\phi$ is shown in the appendix A. For future analysis of GW data, it can be trivially included in the Bayesian analysis framework. We choose the unlensed signal model given as
$h_{x}(f;m_{1},m_{2},d_{l},\iota,s_{1z},s_{2z})=\sum_{i=+,\times}F_{x}^{i}h_{x}^{i}(f;m_{1},m_{2},d_{l},\iota,s_{1z},s_{2z}),$ (7.2)
@@
where $i$ denotes the polarization and $x$ denotes the detector and $F^{+}$ and $F^{\times}$ are the antenna patterns associated with the plus and cross polarization. Here $m_{1}$, $m_{2}$ are the masses of the primary and the secondary of the compact binary, $d_{l}$ is the luminosity distance, $\iota$ is the inclination angle (the angle between the orbital angular momentum and the line of sight). $s_{1z}$ and $s_{2z}$ denote the aligned spin components of the primary and the secondary. The lensed waveform in the frequency domain is the product of the amplification factor and the unlensed signal. Therefore, the lensed signal model is
$h^{lensed}_{x}(f;\{\theta\})=\sum_{i=+,\times}F_{x}^{i}\left[A(f;a,b,k,f_{0})h_{x}^{i}(f;m_{1},m_{2},d_{l},\iota,s_{1z},s_{2z})\right],$ (7.3)
@@
here $\{\theta\}\in\{k,b,f_{0},m_{1},m_{2},d_{\rm eff},\iota,s_{1z},s_{2z}\}$. Notably, we absorb the effect strong lensing magnification $a$ into the luminosity distance of the source $d_{l}$ to combine an effective distance parameter $d_{eff}$. Therefore, in the joint source and lens characterization we include source masses, spins, effective luminosity distance and inclination angle along with microlensing modulations.
We formulate the Bayesian parameter estimation using the package BILBY [63], a python-based module extensively used for GW data analysis purposes. We choose uniform priors for the parameters shown in table 1.
@@
Table 1: Prior distributions for various source and microlensing parameters.
@@
The likelihood is a Gaussian and it uses the frequency-domain data and model from all three detectors (in future we can include KAGRA[9] and LIGO-India [64]). The log-likelihood is given by
$\log(L)=-\sum_{x}\left[\frac{(d_{x}(f)-h_{x}^{lensed}(f;\{\theta\}))^{2}}{2S_{x}(f)}+\frac{1}{2}\log(2\pi S_{x}(f))\right],$ (7.4)
@@
where $d_{x}(f)$ is the data and $S_{x}(f)$ is the noise power spectrum density in detector denoted by $x$. We sample the posteriors using nested sampling algorithm DYNESTY [65]. The corner plot consisting of the parameters is shown in figure 10.
Figure 10: The corner plot shows the distribution of source and lensing parameters obtained from a joint parameter estimation technique in presence and in absence of microlensing.
@@
The figure 10 shows the joint parameters of the source and the lens estimated from the data (i) when there is a microlensed GW signal (ii) when there is no microlensed GW signal. The figure has many different aspects to discuss. Firstly, when the microlensing was present in the data, we note that all the parameters are close to well constrained around their injected values, as compared to the other case when the data did not contain a microlensed GW. In the scenario without microlensing, $k$ value inclined towards zero, $b$ also has a median less than 0.1 and $f_{0}$ has peaked around 190 Hz which is higher than $f_{\rm ISCO}$ for these binaries and close to the maximum frequency of emission from this system. The source parameters are well constrained showing almost no deviation in the parameter distributions with microlensing hypothesis. The joint parameter distribution clearly indicates that there is no degeneracy among the source and lens parameters.
On the other hand, we note that the lensing parameters ($b,f_{0},k$) show no degeneracy with the source parameters ($m_{1},m_{2},d_{\rm eff},\iota,s_{1z},s_{2z}$), in particular the oscillation parameter $f_{0}$ shows no correlation with the source masses $m_{1}$, $m_{2}$, which control the chirping nature of the signal. We also find that the perturbation parameter $b$ is non-degenerate with the aligned spins $s_{1z}$ and $s_{2z}$ which often introduce beating pattern like envelopes on the waveform. This is evident from the fact that, lensing oscillations ($f_{0}$) and its strength ($b$) are coming from the properties of the lens whereas the masses and spins of BBH are its intrinsic properties. Source binary and lensing object involve different timescales of evolution that do not match, therefore the effect of each does not replicate the other, therefore such a correlation between the source and lens parameters is not there. However, we observe that since $a$ and $d_{l}$ for a microlensing event cannot be estimated individually, we combined them in the effective distance parameter $d_{\rm eff}\equiv d_{l}/a$, which is well constrained in the range around injected $d_{\rm eff}$. Finally, we also note that, there is a correlation between $b$ and $k$, which was also observed in the figure 9, since the effect of smaller b is equivalent to making b fixed and k larger.
Here we have considered a prior of $f_{0}$ up to $1024$ Hz. In appendix B, we have shown that even with a reduced prior of $f_{0}$ upto $f_{\rm ISCO}$, the results of the joint distributions and the best-fit reconstruction has only been affected marginally for the same noise realization. This indicates that use of a large prior on $f_{0}$ does not adversely impact the results.
8 False Alarm Rate for Microlensing Detection
To understand the significance of a microlensing detection, we have assigned a False Alarm Rate (FAR) with the detections. The false alarms are triggered when there is actually no microlensing signature present in the data, but $\mu$-GLANCE shows some non-zero cross-correlation signal, caused by the randomness of the noise. We have generated detector noise for Hanford, Livingston, and Virgo with aforementioned noise PSDs for 1 year of observation time. We then performed the cross-correlation between the noise segments of $8$ second length of each with a cross-correlation timescale of $0.5$ second. To calculate the FAR (in units of per year), we obtain the residual cross-correlation signal with varying the source and lens parameters and obtain the number of noise cross-correlation points above the certain residual cross-correlation.
In figure 11, we have showed the cumulative distribution function (CDF) of the noise cross-correlation. We have plotted different lines to show the residual cross-correlation strength variation with network match-filtering SNR, by changing the chirp mass of the system. We have also shown the strength of the case we have considered in figure 6, with dashed lines  [^11] Considering the symmetry of the noise-cross correlation distribution, we have considered the absolute values of the $N_{xx^{\prime}}$ points and plotted its negative. We also did the same for the residual cross-correlation $x\rightarrow{-|x|}$. This was performed to get the probability of a false alarm probability of a residual cross-correlation directly.. We observe that with the system of detector frame BBH masses as $(m_{1},m_{2})=(35,25)M_{\odot}$ at $1$ Gpc with lensing parameters $a=1.8$, $b=0.4$, $f_{0}=30$ Hz and $k=0.004$, we obtain a false alarm probability of $4\times 10^{-5}$ with the residual cross-correlation in H-L pair. Similarly for L-V and V-H pair, we obtain a false alarm probability of few times $4\times 10^{-3}$ and $3\times 10^{-2}$. Since the noise distribution is generated with $1$ year of simulated noise, the false alarm probabilities directly correspond to the FAR (in yr-1 units) corresponding to the detection of the microlensing search.
From figure 11 we also note that, with increase in the chirp mass and therefore the SNR of the signal, how the residual cross-correlation corresponds to lower FAR. We observe, above $60$ network-SNR in H-L and above $80M_{\odot}$ network-SNR in L-V and V-H, the FAR drops below $10^{-3}$, making a significant microlensing detection possible. Therefore, the microlensing detection is very probabilistic in high-snr (above 60 network match-filter SNR at least) systems. We suggest to call an event microlensing candidate if the FAR falls below $10^{-3}$ in any of the detector pair. In appendix C, we have showed the effect of different lensing parameters and have estimated the minimum network matched filter SNR required for a low FAR detection ($\leq 10^{-3}$ per year).
@@ 8 False Alarm Rate for Microlensing Detection
Figure 11: The figure shows the false alarm probabilities for the residual cross-correlation in different detector pair. We have plotted the CDF of the noise cross-correlation distribution (gray) along with different lines showing the residual cross-correlation peak strength by varying the chirp mass of the system. The color of the lines represent the network matched filter SNR of the microlensed signal. Residual cross-correlations from the figure 6 is shown with dashed lines. The values of $a$, $b$, $f_{0}$ and $k$ are 1.8, 0.4, 30 Hz and 0.004 respectively.

### v2

@@ 6 Cross-correlation Signal Dependency on GW Source and Lens Parameters
Figure 9: The figure shows the variation of the residual cross-correlation with respect to variation in the matched-filtering SNR of the GW signal by varying the chirp mass (detector frame) and lensing parameters (b, $f_{0}$ and k). The cross-correlation timescale is chosen to be $\tau=0.5s$. This shows that where the residual cross-correlation deviates from the noise cross-correlations significantly (when the SNR is large and also for large b, small k), the detection becomes efficient. However, in the regime, when the deviation is small (when the SNR is small and also for small b, large k), the detection is inefficient. The effect of the $f_{0}$ on the results is not very decisive.
@@
The residual cross-correlation $(D_{xx^{\prime}})$ depends on the performance of the unlensed hypothesis parameter estimation to obtain the best-fit GW strain. The better the strain data matches with the best-fit strain is, the smaller is the residual and the residual cross-correlation. However, it is shown in the section 7 that, GW source parameters are well constrained with a model considering no lensing hypothesis even when there is WO microlensing present. This happens because the source parameters are completely non-degenerate with the lensing template considered in the analysis. Therefore, the residual cross-correlation depends on the strength of the signal, formally given by the matched-filtering SNR, that defines how well can we obtain the best-fit. Thus since the WO microlensing parameters are not captured by the lens parameters in the unlensed hypothesis PE, there exists strong dependency of the residual cross-correlation on the values of the WO microlensing parameters. Therefore, to observe how well can we distinguish a lensing modulation, we have checked the variation of the residual cross-correlation with respect to lensing parameters. In figure 9, we have shown the detectability of the residual cross-correlation with respect to the background noise cross-correlation, when no signal is present in the data.
@@
The shaded regions in figure 9, shows the 1-$\sigma$ bounds on the noise cross-correlation distribution. The lines show the residual cross-correlation value when the signal consists of WO microlensing amplification. We varied the chirp mass to increase the strength of the signal. We fixed the value of $a$ to 1.2 and 1.8, and then varied the parameters of $b$, $f_{0}$ and $k$. The luminosity distance of the source is kept fixed at $1$ Gpc. The values of the WO microlensing parameters are chosen in consistency with the numbers obtained in 3 when we fitted the amplification models of point-mass and SIS with our template in figure 4.
@@
In the first row of the plots, we have varied the value of the WO microlensing amplitude parameter $b$, keeping $f_{0}$ and $k$ fixed at $30$ Hz and $0.001$ respectively. We observe that for $a=1.8$, $b=0.4$ at chirp mass as low as $20M_{\odot}$ the cross-correlation signal at H-L lies 3-$\sigma$ above, however at L-V pair the cross-correlation signal comes above the noise distribution 3-$\sigma$ at chirp mass $40M_{\odot}$, same for V-H is at around $25M_{\odot}$. As the value of b is decreased to a lower value, the minimum chirp mass at which the cross-correlation signal is well distinguishable  [^11] By ”well distinguishable” we mean that, the residual cross-correlation is at least 3-$\sigma$ away from the noise cross-correlation distribution of the BBH shifts towards higher masses. Similarly for $a=1.2$ with $b=0.4$, $f_{0}=30$ Hz, $k=0.001$ we observe that for H-L, L-V, and V-H the cross-correlation signal is distinguishable at $30M_{\odot}$, $50M_{\odot}$ and $35M_{\odot}$ respectively, which is moved towards higher values as the value of b is decreased.
@@
In the second row of the plots, we have varied the value of the WO microlensing oscillation parameter $f_{0}$, keeping $b$ and $k$ fixed at $0.4$ and $0.001$ respectively. Unlike the cross-correlation dependency on $b$, we don’t observe a direct observational effect of $f_{0}$ variation, for any of the detector pairs. For both the cases with $a=1.8$ and $a=1.2$, we observe that the three lines corresponding to $f_{0}=30$ Hz, $f_{0}=60$ Hz, $f_{0}=100$ Hz, the lines are intertwined, not significantly altering the minimum chirp mass at which the cross-correlation signal is distinguishable. We understand that, the value of $b$ is more critical to determining WO microlensing in GW data as compared to $f_{0}$.
@@
In the third row of the plots, we have varied the value of the WO microlensing oscillation converging parameter $k$ (as the GW frequencies move towards geometric-optics lensing regime), keeping $b$ and $f_{0}$ fixed at $0.4$ and $30$ Hz respectively. Quite similar the cross-correlation dependency on $b$, we observe a strong dependency of $k$ variation. For the case with $a=1.8$, we can distinguish the cross-correlation signal for H-L, L-V, and V-H at $35M_{\odot}$, $45M_{\odot}$ and $40M_{\odot}$ respectively for the value of $k=0.008$. The minimum chirp mass is moved towards the lower end as the value of $k$ is decreased. Similarly, with $a=1.2$, we can distinguish the cross-correlation signal for H-L, L-V, and V-H at $40M_{\odot}$, $60M_{\odot}$ and $45M_{\odot}$ respectively for the value of $k=0.008$. We understand that the effect of the transition parameter $k$ is similar to the effect of perturbation parameter $b$, therefore making it a very crucial parameter for the detection of WO microlensed signal.
7 Joint exploration of GW Source Parameters and Lensing Template Parameters
@@
Lensing can affect the inference of the source properties obtainable from the GW signal through existing parameter estimation techniques. Therefore, a correct approach for inferring source parameters, not just includes the source intrinsic and extrinsic parameters, rather it as well consists of the lens-induced characteristic parameters on the GW signal. We perform Bayesian parameter estimation technique in the frequency domain to jointly explore the parameter space occupied by the source and lensing template parameters. To showcase the capability of the technique, we choose a simple model for the lensing amplification factor as shown in equation 10. Here, we have kept the value of $\phi=0$ for the simulation set for the purpose of demonstrating the method. The results including $\phi$ is shown in the appendix A. For future analysis of GW data, it can be trivially included in the Bayesian analysis framework. We choose the unlensed signal model given as
$h_{x}(f;m_{1},m_{2},d_{\rm source},\iota,s_{1z},s_{2z})=\sum_{i=+,\times}F_{x}^{i}h_{x}^{i}(f;m_{1},m_{2},d_{\rm source},\iota,s_{1z},s_{2z}),$ (11)
Figure 10: The corner plot shows the distribution of source and lensing template parameters obtained from a joint parameter estimation technique in presence and in absence of WO microlensing. It shows that for WO microlensing present and WO microlensing absent cases, the source parameter estimations are not very different, although quite expectedly, the lensing parameters are occupying different regions of the parameter space.
@@
where $i$ denotes the polarization and $x$ denotes the detector and $F^{+}$ and $F^{\times}$ are the antenna patterns associated with the plus and cross polarization. Here $m_{1}$, $m_{2}$ are the masses of the primary and the secondary of the compact binary, $d_{\rm source}$ is the luminosity distance to the source, $\iota$ is the inclination angle (the angle between the orbital angular momentum and the line of sight). $s_{1z}$ and $s_{2z}$ denote the aligned spin components of the primary and the secondary. The lensed waveform in the frequency domain is the product of the amplification factor and the unlensed signal. Therefore, the lensed signal model is
[…]
hxi(f;m1,m2,dsource,ι,s1z,s2z)],
@@
here $\{\theta\}\in\{k,b,f_{0},m_{1},m_{2},d_{\rm eff},\iota,s_{1z},s_{2z}\}$. Notably, we absorb the effect strong lensing magnification $a$ into the luminosity distance of the source $d_{\rm source}$ to combine an effective distance parameter $d_{\rm eff}$. Therefore, in the source and lensing template parameters joint characterization we include source masses, spins, effective luminosity distance and inclination angle along with WO microlensing modulations.
We formulate the Bayesian parameter estimation using the package BILBY (Ashton et al., 2019a), a python-based module extensively used for GW data analysis purposes. We use the waveform model IMRPhenomXPHM for both data generation and Bayesian parameter estimation purposes. We highlight that the estimation of source and lensing template parameters is waveform model-dependent and can vary with different waveform prescriptions due to variations in underlying physical assumptions and approximations. However, this waveform systematics effect can be mitigated by choosing a time interval between which different waveform models agree with one another. This is achieved by choosing the maximum frequency cut-off as $f_{\rm isco}$, up to which the signal is at the inspiral epoch, during which the waveform systematics are limited. For the priors, we choose uniform distribution for all parameters as mentioned in table 1.
@@
Table 1: Prior distributions for various source and WO microlensing parameters.
@@
The likelihood is a Gaussian and it uses the frequency-domain data and model from all three detectors (in future we can include KAGRA(Akutsu et al., 2021) and LIGO-India (Iyer et al., 2011)). The log-likelihood is given by
$\log(L)=-\sum_{f}\sum_{x}\left[\frac{(d_{x}(f)-h_{x}^{\rm lensed}(f;\{\theta\}))^{2}}{2S_{x}(f)}+\frac{1}{2}\log(2\pi S_{x}(f))\right],$ (23)
@@
where $d_{x}(f)$ is the data and $S_{x}(f)$ is the noise power spectrum density in detector denoted by $x$. We sample the posteriors using nested sampling algorithm DYNESTY (Speagle, 2020). The corner plot consisting of the parameters is shown in figure 10.
@@
The figure 10 shows the distribution of parameters of the source and the lensing estimated from the data (i) when there is a WO microlensed GW signal (ii) when there is no WO microlensed GW signal. The figure has many different aspects to discuss. Firstly, when the WO microlensing was present in the data, we note that all the parameters are close to well constrained around their injected values, as compared to the other case when the data did not contain a WO microlensed GW. In the scenario without WO microlensing, $k$ value inclined towards zero, $b$ also has a median less than 0.1 and $f_{0}$ has peaked around 190 Hz which is higher than $f_{\rm isco}$ for these binaries and close to the maximum frequency of emission from this system. The source parameters are well constrained showing almost no deviation in the parameter distributions with WO microlensing hypothesis. The joint parameter distribution clearly indicates that there is no degeneracy among the source and lensing template parameters.
@@
Figure 11: The figure shows the false alarm probabilities for the residual cross-correlation in different detector pair. We have plotted the CDF of the noise cross-correlation distribution (gray) along with different lines showing the residual cross-correlation peak strength by varying the chirp mass of the system. The colour of the lines represent the network matched filter SNR of the WO microlensed signal. Residual cross-correlations from the figure 6 is shown with dashed lines. The values of $a$, $b$, $f_{0}$ and $k$ are 1.8, 0.4, 30 Hz and 0.004 respectively.
On the other hand, we note that the lensing parameters ($b,f_{0},k$) show no degeneracy with the source parameters ($m_{1},m_{2},d_{\rm eff},\iota,s_{1z},s_{2z}$), in particular the oscillation parameter $f_{0}$ shows no correlation with the source masses $m_{1}$, $m_{2}$, which control the chirping nature of the signal. We also find that the perturbation parameter $b$ is non-degenerate with the aligned spins $s_{1z}$ and $s_{2z}$ which often introduce beating pattern like envelopes on the waveform. This is evident from the fact that, lensing oscillations ($f_{0}$) and its strength ($b$) are coming from the properties of the lens whereas the masses and spins of BBH are its intrinsic properties. Note that lensing effects cannot change the GW frequency evolution. The chirping nature of GW frequency evolution is solely dependent on the chirp mass of compact binary. Thus the frequency evolution is controlled by the GW source and frequency-dependent modulations are controlled by the lens properties. Therefore, from a lensed GW, the effect of WO microlensing features can be distinguished from the source imprinted properties, therefore such a correlation between the source and lensing parameters can be mitigated. However, we observe that since $a$ and $d_{\rm source}$ for a lensed event cannot be estimated individually, we combined them in the effective distance parameter $d_{\rm eff}\equiv d_{\rm source}/a$, which is well constrained in the range around injected $d_{\rm eff}$. Finally, we also note that, there is a correlation between $b$ and $k$, which was also observed in the figure 9, since the effect of smaller b is equivalent to making b fixed and k larger.
Here we have considered a prior of $f_{0}$ up to $1024$ Hz. In appendix B, we have shown that even with a reduced prior of $f_{0}$ up to the frequency $f_{\rm isco}$, the results of the joint distributions and the best-fit reconstruction has only been affected marginally for the same noise realization. This indicates that use of a large prior on $f_{0}$ does not adversely impact the results.
8 False Alarm Rate for WO microlensing Detection
To understand the significance of a WO microlensing detection, we have assigned a False Alarm Rate (FAR) with the detections. The false alarms are triggered when there is actually no WO microlensing signature present in the data, but $\mu$-GLANCE shows some non-zero cross-correlation signal, caused by the randomness of the noise. We have generated detector noise for Hanford, Livingston, and Virgo with aforementioned noise PSDs for 1 year of observation time. We then performed the cross-correlation between the noise segments of $8$ second length of each with a cross-correlation timescale of $0.5$ second. To calculate the FAR (in units of per year), we obtain the residual cross-correlation signal with varying the source and lens parameters and obtain the number of noise cross-correlation points above the certain residual cross-correlation.
In figure 11, we have showed the cumulative distribution function (CDF) of the noise cross-correlation. We have plotted different lines to show the residual cross-correlation strength variation with network matched-filtering SNR, by changing the chirp mass of the system. We have also shown the strength of the case we have considered in figure 6, with dashed lines  [^12] Considering the symmetry of the noise-cross correlation distribution, we have considered the absolute values of the $N_{xx^{\prime}}$ points and plotted its negative. We also did the same for the residual cross-correlation $x\rightarrow{-|x|}$. This was performed to get the probability of a false alarm probability of a residual cross-correlation directly.. We observe that with the system of detector frame BBH masses as $(m_{1},m_{2})=(35,25)M_{\odot}$ at $1$ Gpc with lensing parameters $a=1.8$, $b=0.4$, $f_{0}=30$ Hz and $k=0.004$, we obtain a false alarm probability of $4\times 10^{-5}$ with the residual cross-correlation in H-L pair. Similarly for L-V and V-H pair, we obtain a false alarm probability of few times $4\times 10^{-3}$ and $3\times 10^{-2}$. Since the noise distribution is generated with $1$ year of simulated noise, the false alarm probabilities directly correspond to the FAR (in yr-1 units) corresponding to the detection of the WO microlensing search.
From figure 11 we also note that, with increase in the chirp mass and therefore the SNR of the signal, how the residual cross-correlation corresponds to lower FAR. We observe, above $60$ network-SNR in H-L and above $80M_{\odot}$ network-SNR in L-V and V-H, the FAR drops below $10^{-3}$ per year, making a robust WO microlensing detection possible. Therefore, the WO microlensing detection is very likely in high-SNR (above 60 network matched-filter SNR at least) systems. We suggest to call an event WO microlensed candidate if the FAR falls below $10^{-3}$ per year in any of the detector pair. In appendix C, we have showed the effect of different lensing parameters and have estimated the minimum network matched filter SNR required for a low FAR detection ($\leq 10^{-3}$ per year).